OpenStudy (anonymous):

Given the equation y - 4 = three fourths(x + 8) in point-slope form, identify the equation of the same line in standard form. -three fourthsx + y = 10 3x - 4y = -40 y = three fourthsx + 12 y = three fourthsx + 10

OpenStudy (anonymous):

@TuringTest

OpenStudy (turingtest):

standard form is\[y=mx+b\]where \(m\) is again the slope, and \(b\) is the y-intercept (not the same \(b\) as the last problem; don't get confused) Just simplify \(y-4=\frac34(x+8)\) and solve it for \(y\) to get your answer

OpenStudy (turingtest):

\[y-4=\frac34(x+8)\]distribute the \(\frac34\) and what do we get?

myininaya (myininaya):

Some text say Ax+By=C is standard form.

OpenStudy (turingtest):

oh dang, I'll let you take it then myin :/

myininaya (myininaya):

Are you sure?

OpenStudy (turingtest):

yeah i want to eat lol

OpenStudy (anonymous):

you are leaving???

myininaya (myininaya):

Well First I would do what turing as said and distribute that 3/4

OpenStudy (anonymous):

ok

OpenStudy (turingtest):

don't worry, I leave you in extremely capable hands @math_help101

OpenStudy (anonymous):

oh ok well bye have fun eating enjoy:) @TuringTest

OpenStudy (anonymous):

ok may you help me with distributing it please @myininaya

myininaya (myininaya):

a(b+c)=ab+ac is distributive property

OpenStudy (anonymous):

ok so how would i distribute 3/4

myininaya (myininaya):

to both terms inside the ( )

myininaya (myininaya):

just like a(b+c)=ab+ac has explained

myininaya (myininaya):